# The orbital evolution of the tidally stripped star and disk-driven stable mass transfer for QPEs in GSN 069

### v4

@@ Abstract
Context. The origin of the quasi-periodic eruptions (QPEs) is possibly mass loss at the periastron of a body moving around the supermassive black hole (SMBH) in a high eccentric orbit. Such a tidally stripped star is expected to radiate gravitational wave thereby leading to shrinkage of the periastron distance, and thus will eventually be disrupted by the SMBH in the previous studies.
Aims. This scenario predicts a gradually increasing mass transfer, contradicting the long term evolution of the observed intensity of the QPEs in GSN 069.
Methods. In this paper, we first revisited the orbital evolution of the stripped star. Then we suggested a model of a tidal stripped WD moving inside an accretion disk for QPEs in GSN 069.
Results. We found the effect of the mass transfer finally dominates the orbital evolution, resulting in the stripped star finally escaping the SMBH rather than being disrupted by it. The drag force by the disk can effectively reduce mass transfer and thus explain the observed long term evolution in the intensity of the QPEs in GSN 069. The disk is likely a fallback disk of the tidal disruption event in GSN 069. Considering the evolution of its accretion rate, the increase in the intensity of the latest eruption can also be explained.
@@ 1 Introduction
A white dwarf (WD) orbiting around a supermassive black hole (SMBH) with high eccentricity may lose mass each time through the periastron. Such a tidally stripped WD is not only expected to emit periodic electromagnetic wave signals (Zalamea et al., 2010), but also the target of space-based gravitational wave detectors such as Laser Interferometer Space Antenna (LISA) (Babak et al., 2017) and TianQin (Luo et al., 2016). The first discovered quasi-periodic eruptions (QPEs) from GSN 069 are suggested to possibly originate from a tidally stripped WD in a high eccentric orbit (Miniutti et al., 2019; Wang et al., 2022; King, 2020).
@@
Due to gravitational wave radiation decreasing the periastron distance, the mass loss of the tidally stripped WD gradually increases until finally the WD is tidally disrupted by the SMBH (Zalamea et al., 2010; Wang et al., 2022). However, the intensity of QPEs in GSN 069 shows no sign of increasing (Miniutti et al., 2023b, a), which is inconsistent with the previous prediction. King (2022) points out that the mass transfer stability requires the angular momentum transfer from the disk to the orbiting star, he suggested that orbital resonance with the disk possibly could accomplish this, but further discussion was lacking.
@@
GSN 069 also has a tidal disruption event(TDE) that has lasted more than a decade (Shu et al., 2018; Sheng et al., 2021). This TDE was recently found to have re-brightened, and while it was brightening, the QPEs disappeared (Miniutti et al., 2023b). When radiation from the TDE faded again, the QPEs reappeared and were stronger than in the last outburst before they disappeared (Miniutti et al., 2023a). Such behavior implies that the evolution of mass transfer may be related to the accretion rate of the accretion disk.
@@
The origin of this TDE is likely the disruption of a common envelope of the binary containing the current orbiting body before it is captured by the SMBH (Wang, 2023). This means that the fallback disk of the TDE is nearly coplanar with the orbit of the orbiting body that produced the QPEs. In this situation, the drag force by the disk may significantly affect the secular orbital evolution of the orbiting body.
@@
In this paper, we first revisited the secular evolution of the tidally stripped star with high eccentricity taking into account the effect of the mass transfer on the orbit, which is ignored in the previous studies. We found that the effect of mass transfer significantly changes the criterion for unstable mass transfer. The mass transfer for WDs is no longer dynamically stable, while it remains dynamically stable for main sequence stars. Furthermore, the effect of the mass transfer dominates the secular orbital evolution, which leads to the fate of the stripped star escaping the SMBH rather than being disrupted by it.
@@
While the origin of QPEs is still debated, there are various other models (Chen et al., 2022; Zhao et al., 2022; Wang et al., 2022; Krolik & Linial, 2022; Lu & Quataert, 2023; Xian et al., 2021; Suková et al., 2021; Linial & Sari, 2023; Linial & Metzger, 2023; Franchini et al., 2023). In this paper, we focus on the stripped WD model in a high eccentric orbit for GSN 069 (King, 2020, 2022). We calculated the secular effect of drag force by the disk on the orbit for GSN 069. We found that the drag force can transfer the angular momentum of the disc to the stripped WD, and then stabilize the mass transfer. This effect is strongly dependent on the accretion rate, which can be estimated from the luminosity of the TDE. The derived change in mass transfer rate by the disk drag has the ability to suppress other effects that increase mass transfer, thus explaining the long-term evolution of the observed intensity of QPEs. The increase in the intensity of the QPEs after the TDE re-brightened can also be explained by this model.
@@
This paper is organized as follows. In Section 2, we revisited the secular orbital evolution of the stripped star under the effect of gravitational wave radiation and mass transfer. In Section 3, we discuss the effect of drag force by the accretion disk on the orbit for QPEs in GSN 069. We made discussions in section 4 and a summary in Section 5.
@@ 2 Evolution with mass transfer and gravitational radiation
We consider a WD of mass $M$ and radius $R$ in a highly eccentric orbit around a SMBH of mass $M_{h}$. When it reaches the periastron, the stripped mass can be obtained by calculating the mass of the matter in the WD outside Roche lobe (Chen et al., 2023)
@@
Here $\beta\equiv r_{t}/r_{p}$ is impact factor, where $r_{t}=R(M_{h}/M)^{1/3}$ is the disrupted radius and $r_{p}$ is the periastron radius. The Chandrasekhar mass $M_{\mathrm{ch}}$ is $1.44M_{\odot}$ and $\beta_{0}$ is about 0.5. Mass transfer occurs only if $\beta$ is greater than $\beta_{0}$. And the WD is disrupted when $\beta$ is greater than 1.
Differentiating it gives
@@
The latter term is usually dominant because $\beta$ is close to $\beta_{0}$ when mass transfer is not extremely high. Setting the mass-radius index $\zeta$ which gives mass-radius relation $R\propto M^{\zeta}$, the $\dot{\beta}$ can be obtained by (King, 2022)
@@
Here $q=M/M_{h}$ is the mass ratio, $J$ is the total orbital angular momentum and $e$ is the orbital eccentricity.
@@ 2.1 Mass transfer
The mass-radius relation of the WD is (Paczynski, 1971)
@@
which is adopted in the derivation of Eq. (1) (Chen et al., 2023). So $\zeta$ for the WD is about -1/3 for low mass, and more negative when mass is close to $M_{\mathrm{ch}}$. And it is always greater than -5/3 so that the first term in Eq. (3) is negative, which can suppress mass transfer. King (2022) claim that this term is dominant on short timescales so that the mass transfer is dynamically stable. However, he ignored the $\dot{e}$ caused by the mass transfer, which is [^1] Wang et al. (2022) had considered the effect of the mass transfer on the orbit too. Their formula is different from this because they mistake the instantaneous mass transfer rate in Sepinsky et al. (2007b) for the secular mass transfer rate thus underestimate this effect. (Sepinsky et al., 2007b)
@@
Here $\langle\dot{J}/J\rangle\approx 0$, because only very little angular momentum is removed by central SMBH via accretion(King, 2022).
So the self-driven mass transfer for the WDs is unstable. Noted that Eq. (5) show that orbit is more eccentric under mass transfer, and the change of the semi-major axis (Sepinsky et al., 2007b),
@@ 2.2 Gravitational radiation
The change of orbit caused by the gravitational radiation is (Peters, 1964)
@@
Here $G$ is the gravitational constant, $M_{t}=M_{h}+M$ is total mass and $c$ is the speed of light. Substituting Eq. (8) and (10) into Eq. (3) yields
@@ 2.3 Orbital Evolution for the WD
The effect of the gravitational radiation is stronger than the effect of the mass transfer when the mass transfer rate is low. As the mass transfer rate gets higher, the effect of mass transfer becomes dominant. While factor $1/(1-e)$ evolves to very high values, the effect of gravitational radiation becomes smaller. Because the periastron $r_{p}=a(1-e)$ is constant under mass transfer, $\langle\dot{r_{p}}\rangle_{\dot{M}}=0$, which make the effect of gravitational radiation proportional to $(1-e)^{1/2}$ for $\dot{a}/a$ and $(1-e)^{3/2}$ for $\dot{J}/J$ and $\dot{e}$ when $e\approx 1$.
The mass transfer rate $|\dot{M}|$ doesn’t grow endlessly. Because $\dot{M}$ is related not only to $\Delta M$ but also to the orbital period $P_{b}$, $\dot{M}=-\Delta M/P_{b}$. And the $a$ grows dramatically at high $e$, as does the $P_{b}$. When the growth rate of $P_{b}$ is greater than the growth rate of $\Delta M$, $|\dot{M}|$ will start to decrease until the WD escapes. And the critical condition can be obtained by
@@
So the WD finally escapes with eccentricity $e_{f}\gtrsim 1$, then the total mass it loses can be obtained from Eq. (5)
@@
Here we show an example of the evolution under mass transfer and gravitational radiation in Fig. 1. We adopt possible parameters of GSN 069: $P_{b}=9\mathrm{\leavevmode\nobreak\ hr}$, $M_{h}=5\times 10^{5}M_{\odot}$ and $M=0.21M_{\odot}$. Then $e=0.968$ can be obtained from Eq. (4) with $\beta\approx\beta_{0}$. The initial $\beta$ is 0.50001. The orbital parameters and the mass and radius of the WD are updated each time the WD passes through the periastron. We use $\langle\dot{r_{p}}\rangle$ and $\langle\dot{e}\rangle$ to evolve orbit, because $\langle\dot{a}/a\rangle$ is too high to properly represent orbital changes in the above method when $1-e$ is very small. Evolution will stop when $e$ is greater than 1.
Figure 1: The evolution of orbit parameters and WD’s mass. The first panel from top to bottom shows the evolution of $1-e$ and orbital period $P_{b}$. The second one is the evolution of the average mass transfer rate and mass of the WD. And the last one represents the evolution of mass loss at periastron. The dashed line represents the critical time that the effect of the mass transfer becomes dominant. For each parameter, the critical times are not the same but very close. Here we use the critical time of $e$ for all panels. The last 10 orbits are plotted as scatter plots and the rest as line plots.
@@
2.4 Orbital Evolution for the main-sequence star
If $\zeta$ of the stripped star is greater than 1/3, such as a main-sequence star, then the mass transfer will stabilize itself, and $\langle\dot{\beta}\rangle$ will evolve towards 0. Once this happens, $\langle\dot{\beta}\rangle=0$ will remain. Since the timescale of the mass transfer evolution is larger than that of the orbital parameter evolution, $\dot{M}$ can be able to adjust itself rapidly to balance with the effects of gravitational radiation in the secular evolution. That is, it always satisfies
@@ 2.4 Orbital Evolution for the main-sequence star
Figure 2: Comparison of $\langle\dot{a}\rangle$ and $\langle\dot{e}\rangle$ caused by the mass transfer and gravitational radiation with $\langle\dot{\beta}\rangle=0$. Here $\eta$ is taken as 2/3.
While these two effects are balanced on $\langle\dot{\beta}\rangle$, they are not balanced on $\langle\dot{a}\rangle$ and $\langle\dot{e}\rangle$. As shown in Fig. 2, the effect of the mass transfer is always dominant when $\langle\dot{\beta}\rangle=0$, but it is of the same magnitude as GW. Thus, the evolution of a stripped main-sequence is similar to that of a WD, but on a much longer timescale, the gravitational radiation timescale, because $\dot{M}$ is driven by gravitational radiation.
Figure 3: Similar as Fig. 1, but for main-sequence star. The last panel from the top also shows the evolution of $\langle\dot{\beta}\rangle$ of gravitational radiation and mass transfer.
We show an example in Fig. 3. The initial orbital parameters are $M_{h}=5\times 10^{5}M_{\odot}$, $M=0.1M_{\odot}$, $\beta=0.50004$ and $P_{b}=40\mathrm{\leavevmode\nobreak\ hr}$. The radius is obtained by $R=1.06(M/M_{\odot})^{0.945}R_{\odot}$ (Demircan & Kahraman, 1991). Then the eccentricity can be obtained $e=0.912$.
As shown in Fig. 3, the effect of mass transfer is initially weak and $|\dot{M}|$ increase rapidly driven by gravitational radiation, and then reaching balance $\langle\dot{\beta}\rangle=0$. At this point, the evolution of the orbit is dominated by the effects of mass transfer, with $r_{p}$ remaining almost constant while $e$ grows. Since $\langle\dot{\beta}/\beta\rangle_{GW}$ scales as $(1-e)^{3/2}$, $|\dot{M}|$ coupled to it will keep decreasing. Eventually, like WDs, main sequence stars escape from SMBH, but on much longer timescales.
There is also a special case where, if the initial $|\dot{M}|$ and $e$ are both high, $e$ may evolve to greater than 1 before the $|\dot{M}|$ has had time to be reduced to balance state. In addition, in the final stages of evolution, when the orbital period is growing rapidly, there is a possibility that the star may be perturbed by other gravitational sources away from the black hole and thus escape early.

### v5

# Orbital evolution of a tidally stripped star and disk-driven stable mass transfer for QPEs in GSN 069

@@ Abstract
Context. A plausible origin for the quasi-periodic eruptions (QPEs) could be mass loss at the periastron of a body moving around the supermassive black hole (SMBH) on a highly eccentric orbit. This kind of tidally stripped star is expected to radiate gravitational waves, thereby leading to a shrinkage of the periastron distance. As a result, it will eventually be disrupted by the SMBH, as suggested by previous studies.
Aims. This scenario predicts a gradually increasing mass transfer, contradicting the long-term evolution of the observed intensity of QPEs in GSN 069.
Methods. In this work, we revisit the orbital evolution of the stripped star and we propose a model of a tidally stripped WD moving inside an accretion disk for QPEs, aimed at characterizing GSN 069.
Results. We found the effect of the mass transfer ultimately dominates the orbital evolution, resulting in the stripped star finally escaping the SMBH, rather than being disrupted by it. The drag force induced by the disk may effectively reduce the mass transfer and could thus explain the observed long-term evolution in the intensity of the QPEs in GSN 069. The disk is likely a fallback disk of the tidal disruption event in GSN 069. Considering the evolution of its accretion rate, this scenario could also explain the increase in the intensity of the latest eruption.
@@ 1 Introduction
A white dwarf (WD) orbiting around a supermassive black hole (SMBH) with a high eccentricity may lose mass each time via the periastron. Such a tidally stripped WD is not only expected to emit periodic electromagnetic wave signals (Zalamea et al., 2010), but it is also a target for space-based gravitational wave detectors, such as Laser Interferometer Space Antenna (LISA, Babak et al. 2017) and TianQin (Luo et al., 2016). The first discovered quasi-periodic eruptions (QPEs) from GSN 069 are suggested to have possibly originated from a tidally stripped WD on a highly eccentric orbit (Miniutti et al., 2019; Wang et al., 2022; King, 2020).
@@
Due to gravitational wave radiation decreasing the periastron distance, the mass loss of the tidally stripped WD gradually increases until, finally, the WD is tidally disrupted by the SMBH (Zalamea et al., 2010; Wang et al., 2022). However, the intensity of QPEs in GSN 069 shows no sign of increasing (Miniutti et al., 2023b, a), which is inconsistent with the cited prediction. King (2022) pointed out that the mass transfer stability requires the angular momentum transfer from the disk to the orbiting star. He suggested that an orbital resonance with the disk possibly could accomplish this, but any further discussion is still lacking.
@@
GSN 069 also has a tidal disruption event (TDE) that has lasted more than a decade (Shu et al., 2018; Sheng et al., 2021). This TDE was recently found to have re-brightened and, while it was brightening, the QPEs disappeared (Miniutti et al., 2023b). When radiation from the TDE faded again, the QPEs reappeared, and one of two detected QPEs was stronger than in the last outburst before they disappeared (Miniutti et al., 2023a). Such a behavior implies that the evolution of mass transfer may be related to the accretion rate of the accretion disk.
@@
The origin of this TDE is likely caused by the disruption of a common envelope of the binary containing the current orbiting body before it is captured by the SMBH (Wang, 2024). This means that the fallback disk of the TDE is nearly coplanar with the orbit of the orbiting body that produced the QPEs. In this situation, the drag force by the disk may significantly affect the secular orbital evolution of the orbiting body.
@@
In this paper, we first revisit the secular evolution of the tidally stripped star with high eccentricity, taking into account the effect of the mass transfer on the orbit, which has been ignored in the previous studies. We found that the effect of mass transfer significantly changes the criterion for unstable mass transfer. The mass transfer for WDs is no longer dynamically stable, while it remains dynamically stable for main sequence stars. Furthermore, the effect of the mass transfer dominates the secular orbital evolution, which leads to the fate of the stripped star escaping the SMBH, rather than being disrupted by it.
@@
While the origin of QPEs is still a matter of debate, there are various other models to consider in this context (Chen et al., 2022; Zhao et al., 2022; Wang et al., 2022; Krolik & Linial, 2022; Lu & Quataert, 2023; Xian et al., 2021; Suková et al., 2021; Linial & Sari, 2023; Linial & Metzger, 2023; Franchini et al., 2023). In this paper, we focus on the stripped WD model on a highly eccentric orbit for GSN 069 (King, 2020, 2022). We calculated the secular effect of drag force by the disk on the orbit for GSN 069. We found that the drag force can transfer the angular momentum of the disc to the stripped WD, and then stabilize the mass transfer. This effect is strongly dependent on the accretion rate, which can be estimated from the luminosity of the TDE. The derived change in the mass transfer rate by the disk drag features the ability to suppress other effects that can increase mass transfer, thus explaining the long-term evolution of the observed intensity of QPEs. The increase in the intensity of the partial QPEs after the TDE re-brightened may also be explained by this model.
@@
This paper is organized as follows. In Section 2, we revisit the secular orbital evolution of the stripped star under the effect of gravitational wave radiation and mass transfer. In Section 3, we discuss the effect of drag force by the accretion disk on the orbit for QPEs in GSN 069. We present our discussion in Section 4 and a summary in Section 5.
@@ 2 Evolution with mass transfer and gravitational radiation
We consider a WD of mass $M$ and radius $R$ on a highly eccentric orbit around a SMBH of mass $M_{h}$. When it reaches the periastron, the stripped mass can be obtained by calculating the mass of the matter in the WD outside Roche lobe (Chen et al., 2023):
@@
Here, $\beta\equiv r_{t}/r_{p}$ is impact factor, where $r_{t}=R(M_{h}/M)^{1/3}$ is the disrupted radius and $r_{p}$ is the periastron radius. The Chandrasekhar mass $M_{\mathrm{ch}}$ is $1.44M_{\odot}$ and $\beta_{0}$ is about 0.5. Mass transfer occurs only if $\beta$ is greater than $\beta_{0}$. And the WD is disrupted when $\beta$ is greater than 1.
Differentiating it then gives:
@@
The latter term is usually dominant because $\beta$ is close to $\beta_{0}$ when there is not a great amount of mass transfer. Setting the mass-radius index, $\zeta,$ which gives mass-radius relation of $R\propto M^{\zeta}$, then $\dot{\beta}$ can be obtained via (King, 2022):
@@
Here, $q=M/M_{h}$ is the mass ratio, $J$ is the total orbital angular momentum and $e$ is the orbital eccentricity.
@@ 2.1 Mass transfer
The mass-radius relation of the WD is (Paczynski, 1971):
@@
which is adopted in the Eq. (1) derivation (Chen et al., 2023). So, $\zeta$ for the WD is about -1/3 for a low mass and becomes even more negative when mass is close to $M_{\mathrm{ch}}$. Furthermore, it is always greater than -5/3, so that the first term in Eq. (3) is negative, which can suppress mass transfer. King (2022) claimed that this term is dominant on short timescales, so the mass transfer is dynamically stable. However, he ignored the $\dot{e}$ caused by the mass transfer, which is [^1] Wang et al. (2022) considered the effect of the mass transfer on the orbit as well. Their formula is different from this because they mistake the instantaneous mass transfer rate in Sepinsky et al. (2007b) for the secular mass transfer rate, thus underestimating this effect. (Sepinsky et al., 2007b):
@@
Here, $\langle\dot{J}/J\rangle\approx 0$, because only very little angular momentum is removed by central SMBH via accretion(King, 2022).
So the self-driven mass transfer for the WDs is unstable. Eq. (5) shows the orbit is more eccentric under mass transfer and the change in the semi-major axis (Sepinsky et al., 2007b):
@@ 2.2 Gravitational radiation
The change in orbit caused by the gravitational radiation is (Peters, 1964)
@@
Here, $G$ is the gravitational constant, $M_{t}=M_{h}+M$ is total mass, and $c$ is the speed of light. Substituting Eq. (8) and (10) into Eq. (3) yields
@@ 2.3 Orbital evolution for the WD
The effect of the gravitational radiation is stronger than the effect of the mass transfer when the mass transfer rate is low. As the mass transfer rate gets higher, the effect of mass transfer becomes dominant. While factor $1/(1-e)$ evolves to very high values, the effect of gravitational radiation becomes smaller. Because the periastron $r_{p}=a(1-e)$ is constant under mass transfer, $\langle\dot{r_{p}}\rangle_{\dot{M}}=0$, this makes the effect of gravitational radiation proportional to $(1-e)^{1/2}$ for $\dot{a}/a$ and $(1-e)^{3/2}$ for $\dot{J}/J$ and $\dot{e}$ when $e\approx 1$.
The mass transfer rate $|\dot{M}|$ doesn’t grow endlessly. Because $\dot{M}$ is related not only to $\Delta M$ but also to the orbital period $P_{b}$, $\dot{M}=-\Delta M/P_{b}$. And then $a$ grows dramatically at high $e$, as does $P_{b}$. When the growth rate of $P_{b}$ is greater than the growth rate of $\Delta M$, $|\dot{M}|$ will start to decrease until the WD escapes. And the critical condition can be obtained by:
@@
So, the WD finally escapes with eccentricity $e_{f}\gtrsim 1$, then the total mass it loses can be obtained from Eq. (5):
@@
Here, we show an example of the evolution under mass transfer and gravitational radiation in Fig. 1. We adopted plausible parameters for GSN 069: $P_{b}=9\mathrm{\leavevmode\nobreak\ hr}$, $M_{h}=5\times 10^{5}M_{\odot}$, and $M=0.21M_{\odot}$. Then, $e=0.968$ can be obtained from Eq. (4) with $\beta\approx\beta_{0}$. The initial $\beta$ is 0.50001. The orbital parameters and the mass and radius of the WD are updated each time the WD passes through the periastron. We use $\langle\dot{r_{p}}\rangle$ and $\langle\dot{e}\rangle$ to evolve orbit, because $\langle\dot{a}/a\rangle$ is too high to properly represent orbital changes in the above method, when $1-e$ is very small. The evolution will stop when $e$ is greater than 1.
Figure 1: Evolution of orbit parameters and WD’s mass. The first panel from top to bottom shows the evolution of $1-e$ and orbital period $P_{b}$. The second one is the evolution of the average mass transfer rate and mass of the WD. And the last one represents the evolution of mass loss at periastron. The dashed line represents the critical time that the effect of the mass transfer becomes dominant. For each parameter, the critical times are not the same but very close. Here we use the critical time of $e$ for all panels. The last ten orbits are plotted as scatter plots and the rest as line plots.
@@
2.4 Orbital evolution for the main sequence star
If $\zeta$ of the stripped star is greater than 1/3, such as a main sequence star, then the mass transfer will stabilize and $\langle\dot{\beta}\rangle$ will evolve towards 0. Once this happens, $\langle\dot{\beta}\rangle=0$ will remain. Since the timescale of the mass transfer evolution is larger than that of the orbital parameter evolution, $\dot{M}$ should be able to adjust itself rapidly to balance with the effects of gravitational radiation in the secular evolution. That is, it always satisfies:
@@ 2.4 Orbital evolution for the main sequence star
Figure 2: Comparison of $\langle\dot{a}\rangle$ and $\langle\dot{e}\rangle$ caused by the mass transfer and gravitational radiation with $\langle\dot{\beta}\rangle=0$. Here, $\eta$ is taken as 2/3.
While these two effects are balanced on $\langle\dot{\beta}\rangle$, they are not balanced on $\langle\dot{a}\rangle$ and $\langle\dot{e}\rangle$. As shown in Fig. 2, the effect of the mass transfer is always dominant when $\langle\dot{\beta}\rangle=0$, but it is of the same magnitude as GW. Thus, the evolution of a stripped main sequence star is similar to that of a WD, but on a much longer timescale: the gravitational radiation timescale. This is because $\dot{M}$ is driven by gravitational radiation.
Figure 3: Similar as Fig. 1, but for a main sequence star. The last panel from the top also shows the evolution of $\langle\dot{\beta}\rangle$ of gravitational radiation and mass transfer.
We show an example of this in Fig. 3, where the initial orbital parameters are $M_{h}=5\times 10^{5}M_{\odot}$, $M=0.1M_{\odot}$, $\beta=0.50004,$ and $P_{b}=40\mathrm{\leavevmode\nobreak\ hr}$. The radius is obtained via $R=1.06(M/M_{\odot})^{0.945}R_{\odot}$ (Demircan & Kahraman, 1991). Then, the eccentricity can be obtained $e=0.912$.
As shown in Fig. 3, the effect of mass transfer is initially weak and $|\dot{M}|$ increase rapidly driven by gravitational radiation and then reaching a balance at $\langle\dot{\beta}\rangle=0$. At this point, the evolution of the orbit is dominated by the effects of mass transfer, with $r_{p}$ remaining almost constant, while $e$ grows. Since $\langle\dot{\beta}/\beta\rangle_{GW}$ scales as $(1-e)^{3/2}$, $|\dot{M}|$ coupled to it will keep decreasing. Eventually, as in the case of WDs, main sequence stars escape from SMBH, but on much longer timescales.
There is also a special case where if the initial $|\dot{M}|$ and $e$ are both high, $e$ may evolve to greater than 1 before $|\dot{M}|$ has had time to be reduced to balance state. In addition, in the final stages of evolution, when the orbital period is growing rapidly, there is a possibility that the star may be perturbed by other gravitational sources away from the black hole and, thus, it could escape early.